# Hierarchical Clustering for Conditional Diffusion in Image Generation

### v1

@@ Abstract
Finding clusters of data points with similar characteristics and generating new cluster-specific samples can significantly enhance our understanding of complex data distributions. While clustering has been widely explored using Variational Autoencoders, these models often lack generation quality in real-world datasets. This paper addresses this gap by introducing TreeDiffusion, a deep generative model that conditions Diffusion Models on hierarchical clusters to obtain high-quality, cluster-specific generations. The proposed pipeline consists of two steps: a VAE-based clustering model that learns the hierarchical structure of the data, and a conditional diffusion model that generates realistic images for each cluster. We propose this two-stage process to ensure that the generated samples remain representative of their respective clusters and enhance image fidelity to the level of diffusion models. A key strength of our method is its ability to create images for each cluster, providing better visualization of the learned representations by the clustering model, as demonstrated through qualitative results. This method effectively addresses the generative limitations of VAE-based approaches while preserving their clustering performance. Empirically, we demonstrate that conditioning diffusion models on hierarchical clusters significantly enhances generative performance, thereby advancing the state of generative clustering models.
@@ 1 Introduction
Generative modeling and clustering are two fundamental yet distinct tasks in machine learning. Generative modeling focuses on approximating the underlying data distribution, enabling the generation of new samples (Kingma & Welling, 2014; Goodfellow et al., 2014). Clustering, on the other hand, seeks to uncover meaningful and interpretable structures within data through the unsupervised detection of intrinsic relationships and dependencies (Ward Jr, 1963; Ezugwu et al., 2022), facilitating better data visualization and interpretation. TreeVAE (Manduchi et al., 2023) was recently proposed to bridge these two research directions by integrating hierarchical dependencies into a deep latent variable model. While TreeVAE is effective at hierarchical clustering, it falls short in generating high-quality images. Like other VAE-based models, TreeVAE faces common issues such as producing blurry outputs (Bredell et al., 2023). In contrast, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have recently gained prominence for their superior image generation capabilities, progressively refining noisy inputs to produce sharp, realistic images.
Our work bridges this gap by introducing a second-stage diffusion model conditioned on cluster-specific representations learned by TreeVAE. The proposed framework, TreeDiffusion, combines the strengths of both models to generate high-quality, cluster-specific images, achieving strong performance in both clustering and image generation. The generative process begins by sampling the root embedding of a latent tree, which is learned during training. From there, the sample is propagated from the root to the leaf by (a) sampling a path through the tree and (b) applying a sequence of stochastic transformations to the root embedding along the chosen hierarchical path. Subsequently, the diffusion model leverages the hierarchical information by conditioning its reverse diffusion process on the sampled path representation of the latent tree. A key strength of TreeDiffusion is its ability to generate images for each cluster, providing enhanced visualization of the learned representations, as demonstrated by our qualitative results. The method produces leaf-specific images that share common general properties but differ by cluster-specific features, as encoded in the latent hierarchy. This approach overcomes the generative limitations of VAE-based clustering models like TreeVAE while preserving their clustering performance.
Our key contributions include: (i) a unified framework that integrates hierarchical clustering into diffusion models, and (ii) a novel mechanism for controlling image synthesis. We demonstrate that our approach (a) surpasses the generative limitations of VAE-based clustering models, and (b) produces samples that are both more representative of their respective clusters and closer to the true data distribution.
@@ 2 Related Work
Variational Approaches for Hierarchical Clustering
@@
Since their introduction, Variational Autoencoders (Kingma & Welling, 2014, VAEs) have been often employed for clustering tasks, as they are particularly effective in learning structured latent representations of data (Jiang et al., 2017). Goyal et al. (2017), for example, integrates hierarchical Bayesian non-parametric priors to the latent space of VAEs by applying the nested Chinese Restaurant Processes to cluster the data based on infinitely deep and branching trees. Additionally, hierarchical clustering has been achieved through models such as DeepECT (Mautz et al., 2020) and TreeVAE (Manduchi et al., 2023), both of which grow and learn hierarchical representations during training. While DeepECT aggregates data into a hierarchical tree in a single shared latent space, TreeVAE learns a tree structure posterior distribution of latent stochastic variables. That is, TreeVAE models the data distribution by learning an optimal tree structure of latent variables, resulting in latent embeddings that are automatically organized into a hierarchy, mimicking the hierarchical clustering process. Single-cell TreeVAE (Vandenhirtz et al., 2024, scTree) further extends TreeVAE to cluster single-cell RNA sequencing data by integrating batch correction, facilitating biologically plausible hierarchical structures. Although the aforementioned works have proven effective for clustering, their generative capabilities often fall short, with few offering quantitative or qualitative evaluations of their generative model performance.
Diffusion Models
Diffusion models have largely become state-of-the-art for image generation tasks (Sohl-Dickstein et al., 2015; Ho et al., 2020). Nichol & Dhariwal (2021) and Dhariwal & Nichol (2021) introduced enhancements to the architecture and training procedures of the Denoising Diffusion Probabilistic Model (DDPM). Meanwhile, other works, such as Song et al. (2020) and Salimans & Ho (2022), have made significant strides in reducing the sampling times for diffusion models, addressing one of their primary drawbacks. Song et al. (2023) further introduced consistency models, a new family of models that generate high-quality samples by directly mapping noise to data, enabling fast one-step generation while still allowing multistep sampling to balance computation and sample quality. Moreover, Song & Ermon (2019) proposed an alternative formulation of diffusion modeling through their core-based generative model, known as the noise conditional score network (NCSN). Finally, recent work has further pushed the boundaries of diffusion models by relocating the diffusion process to the latent spaces of autoencoders, as demonstrated in works like LSGM (Vahdat et al., 2021) and Stable Diffusion (Rombach et al., 2022).
@@
One drawback of diffusion models is that their latent variables lack interpretability compared to the latent spaces of VAEs. To leverage the strengths of both approaches, researchers have begun developing architectures that combine the more interpretable latent spaces of VAEs with the advanced generative capabilities of diffusion models. Notable examples include DiffuseVAE (Pandey et al., 2022), Diffusion Autoencoders (Preechakul et al., 2022), and InfoDiffusion (Wang et al., 2023).
Connecting Diffusion with Clustering
@@
The research most closely related to our work focuses on using clustering as conditioning signals for diffusion models to enhance their generative quality. For instance, Adaloglou et al. (2024) propose an approach that utilizes cluster assignments from k-means or TEMI clustering (Adaloglou et al., 2023). Similarly, Hu et al. (2023) introduces a framework that employs the k-means clustering algorithm as an annotation function, generating self-annotated image-level, box-level, and pixel-level guidance signals. Both studies demonstrate the benefits of conditioning on clustering information to improve generative performance without going into the specifics of clustering performance itself. In contrast, our research further investigates which types of clustering information are most beneficial for the model, employing learned latent cluster representations alongside cluster assignments for conditioning. Related to conditioning on clusters, both kNN-Diffusion (Sheynin et al., 2023) and Retrieval-Augmented Diffusion Models (Blattmann et al., 2022) utilize nearest neighbor retrieval to condition generative models on similar embeddings, minimizing the need for large parametric models and paired datasets in tasks like text-to-image synthesis. Diffusion models have also been applied in incomplete multiview clustering to generate missing views to improve clustering performance, as demonstrated by (Wen et al., 2024; 2020). On a different note, (Wang et al., 2024) shows that training diffusion models is equivalent to solving a subspace clustering problem, explaining their ability to learn image distributions with few samples. Additionally, Palumbo et al. (2023) employ post-hoc diffusion models to enhance the generation quality of their multimodal clustering models. However, to the best of our knowledge, there is currently no diffusion model that leverages the hierarchical structure of the data to enhance the interpretability and generative performance of generative clustering models.
@@ 3 Method
We propose TreeDiffusion [^1] The code is publicly available at https://github.com/JoGo175/TreeDiffusion, a two-stage framework consisting of a first-stage VAE-based generative hierarchical clustering model, followed by a second-stage hierarchy-conditional diffusion model. This novel combination of VAEs and diffusion models extends the DiffuseVAE framework introduced by Pandey et al. (2022) to hierarchical clustering. It enables cluster-guided diffusion in unsupervised settings, as opposed to classifier-guided diffusion for labeled data, as introduced by Dhariwal & Nichol (2021). In our framework, TreeVAE (Manduchi et al., 2023) serves as the clustering model, encoding hierarchical clusters within its latent tree structure, where the leaves represent the clusters. A denoising diffusion implicit model (DDIM) (Song et al., 2020), conditioned on the TreeVAE leaves, utilizes these leaf representations to generate improved cluster-conditional samples. Figure 1 illustrates the workflow of TreeDiffusion.
Figure 1: Schematic overview of the TreeDiffusion framework: TreeVAE encodes data into hierarchical latent variables. A path is sampled through the tree, ending at a leaf, with the leaf embedding decoded to generate a reconstruction. The diffusion model leverages the cluster information from TreeVAE by conditioning its reverse process on the sampled leaf and the path embeddings, producing a sharper cluster-specific version of the image.
@@ 3.1 Hierarchical Clustering with TreeVAE
The first part of TreeDiffusion involves an adapted version of the Tree Variational Autoencoder (TreeVAE) by Manduchi et al. (2023). TreeVAE is a generative model that learns to hierarchically separate data into clusters via a latent tree structure. During training, the model dynamically grows a binary tree structure of stochastic variables, ${\mathcal{T}}$. The process begins with a tree composed of a root node and two child nodes and it optimizes the corresponding ELBO over a fixed number of epochs. Afterward, the tree expands by adding two child nodes to an existing leaf node, prioritizing nodes with the highest assigned sample count to promote balanced leaves. This expansion continues iteratively, training only the subtree formed by the new leaves while freezing the rest of the model. This process repeats until the tree reaches a predefined depth or leaf count, alternating between optimizing model parameters and expanding the tree structure.
Let the set ${\mathbb{V}}$ represent the nodes of the tree. Each node corresponds to a stochastic latent variable, denoted as ${\mathbf{z}}_{0},\dots,{\mathbf{z}}_{V}$. Each latent variable follows a Gaussian distribution, whose parameters depend on their parent nodes through neural networks called transformations. The set of leaves ${\mathbb{L}}$, where ${\mathbb{L}}\subset{\mathbb{V}}$, represents the clusters present in the data. Starting from the root node, ${\mathbf{z}}_{0}$, a given sample traverses the tree to a leaf node, ${\mathbf{z}}_{l}$, in a probabilistic manner. The probabilities of moving to the left or right child at each internal node are determined by neural networks termed routers. These decisions, denoted by $c_{i}$ for each non-leaf node $i$, follow a Bernoulli distribution, where $c_{i}=0$ indicates the selection of the left child. The path $\mathcal{P}_{l}$ refers to the sequence of nodes from the root to a leaf $l$. Thus, the latent tree encodes a sample-specific probability distribution of paths. Each leaf embedding, ${\mathbf{z}}_{l}$ for $l\in{\mathbb{L}}$, represents the learned data representations, and leaf-specific decoders use these embeddings to reconstruct or generate new cluster-specific images, i.e. given a dataset ${\bm{X}}$, TreeVAE reconstructs $\hat{{\bm{X}}}=\{\hat{{\bm{X}}}^{(l)}\mid l\in{\mathbb{L}}\}$. The generative model (1) and inference model (2) of TreeVAE are defined as follows:
@@
$q\left(\bm{z}_{\mathcal{P}_{l}},\mathcal{P}_{l}\mid\bm{x}\right)$ $=q\left(\bm{z}_{0}\mid\bm{x}\right)\prod_{i\in\mathcal{P}_{l}\backslash\{0\}}q\left(\mathrm{c}_{pa(i)\rightarrow i}\mid\bm{x}\right)q\left(\bm{z}_{i}\mid\bm{z}_{pa(i)}\right)$ (2)
The objective of the model is to maximize the evidence lower bound (ELBO), which consists of two main components: the reconstruction term $\mathcal{L}_{rec}$ and the KL divergence term, which is broken down into contributions from the root node, internal nodes, and decision probabilities:
@@
In this work, we modify the architectural design of TreeVAE, which originally uses an encoder to project images into flattened representations and relies on MLP layers for subsequent processing. Instead, we utilize convolutional layers throughout the model, which leverage lower-dimensional, multi-channel representations, thereby avoiding flattening the representations. Additionally, we incorporate residual connections to enhance the training stability and model performance. These modifications aim to preserve spatial information and facilitate more efficient learning, making the model particularly effective for image data. Nevertheless, it is important to note that this model encounters the common VAE issue of producing blurry image generations (Bredell et al., 2023). Despite this limitation, the reconstructed images and learned hierarchical clustering still offer meaningful representations of the data, which are used in the second stage of the proposed TreeDiffusion framework.
3.2 Cluster-Conditioned Diffusion
The second part of TreeDiffusion incorporates a conditional diffusion model. Based on the learned latent tree, the first-stage TreeVAE generates hierarchical clusters, which guide the second-stage diffusion process. The diffusion process involves two processes: forward noising and reverse denoising. We assume the same forward process as in standard Denoising Diffusion Probabilistic Models (Ho et al., 2020, DDPM), which gradually introduces noise to the data $\bm{x}_{0}$ over $T$ steps. The intermediate states, $\bm{x}_{t}$ for $t=1,\dots,T$, follow a trajectory determined by a noise schedule ${\beta_{1},\dots,\beta_{T}}$ that controls the rate of data degradation:
@@ 3.2 Cluster-Conditioned Diffusion
$q\left(\bm{x}_{t}\mid\bm{x}_{0}\right)$ $=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},\left(1-\overline{\alpha}_{t}\right)\bm{I}\right),\text{ where }\alpha_{t}=\left(1-\beta_{t}\right)\text{ and }\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}.$ (6)
For the reverse process, we expand the DiffuseVAE framework by Pandey et al. (2022), where a VAE generates the initial, typically blurred images, and a conditional diffusion model refines them to produce sharper, higher-quality outputs. However, instead of using VAE’s reconstructions, our model incorporates both TreeVAE’s hierarchical representations, denoted as ${\bm{c}}_{l}$, and the leaf-specific reconstructions, $\hat{{\bm{x}}}_{0}^{(l)}$ in the reverse process of the DDPM. The tree leaf, which represents the chosen cluster, is selected by sampling from the TreeVAE path probabilities. The leaf-specific reconstructions are concatenated with the reverse process representation $\bm{x}_{t}$ to obtain $\bm{x}_{t-1}$ at each time step $t$. The hierarchical representations, on the other hand, are used as the conditioning input of the U-Net decoder (Nichol & Dhariwal, 2021). The reverse process can be summarized as follows:
$\begin{split}&l\sim p(l|{\bm{x}}_{0}),\\ &p_{\psi}\!\!\left({\bm{x}}_{0:T}\!\mid\!\hat{{\bm{x}}}_{0}^{(l)},{\bm{c}}_{l}\right)=p\left({\bm{x}}_{T}\right)\prod_{t=1}^{T}p_{\psi}\!\!\left({\bm{x}}_{t-1}\!\mid\!{\bm{x}}_{t},\hat{{\bm{x}}}_{0}^{(l)},{\bm{c}}_{l}\right)\!,\\ \end{split}$ (7)
@@
where $\psi$ denotes the parameters of the denoising model and $p(l|{\bm{x}}_{0})$ is the probability that the sample ${\bm{x}}_{0}$ is assigned to leaf $l$. This method ensures that leaves with smaller assignment probabilities are considered, encouraging the diffusion model to perform effectively across all leaves. Consequently, our approach addresses the distinct clusters inherent to TreeVAE, allowing the model to adapt to different clusters and encouraging cluster-specific refinements in the images. This guidance in the image generation process assists the denoising model in learning cluster-specific image reconstructions. For the cluster-specific conditioning information ${\bm{c}}_{l}$, we consider:
${\bm{c}}_{l}=\left\{\begin{array}[]{ll}l&\text{leaf assignment},\\ \bm{z}_{l}&\text{leaf embedding},\\ \bm{z}_{\mathcal{P}_{l}}&\text{set of latent embeddings from root to leaf},\end{array}\right.$ (8)
where the latent embeddings, $\bm{z}_{l}$, are transformed by MLPs and, in the case of $\bm{z}_{\mathcal{P}_{l}}$, added together before giving them as conditioning signals to the U-Net. For the experiments in Section 4.1 and Section 4.2 we use ${\bm{c}}_{l}=\bm{z}_{\mathcal{P}_{l}}$, as we empirically find that it improves upon generative performance, as we show in Section 4.3. To speed up inference, we opt for the DDIM sampling procedures (Song et al., 2020) instead of the standard DDPM (Ho et al., 2020). DDIMs significantly accelerate inference by utilizing only a subset of denoising steps rather than processing all available steps.
By employing a two-stage training strategy, where the conditional diffusion model is trained using a pre-trained TreeVAE model, TreeDiffusion preserves the hierarchical clustering performance of TreeVAE. Hence, we are able to combine the effective clustering of TreeVAE with the superior image generation capabilities of diffusion models.

### v2

# TreeDiffusion: Hierarchical Generative Clustering for Conditional Diffusion

@@ Abstract
Generative modeling and clustering are conventionally distinct tasks in machine learning. Variational Autoencoders (VAEs) have been widely explored for their ability to integrate both, providing a framework for generative clustering. However, while VAEs can learn meaningful cluster representations in latent space, they often struggle to generate high-quality samples. This paper addresses this problem by introducing TreeDiffusion, a deep generative model that conditions diffusion models on learned latent hierarchical cluster representations from a VAE to obtain high-quality, cluster-specific generations. Our approach consists of two steps: first, a VAE-based clustering model learns a hierarchical latent representation of the data. Second, a cluster-aware diffusion model generates realistic images conditioned on the learned hierarchical structure. We systematically compare the generative capabilities of our approach with those of alternative conditioning strategies. Empirically, we demonstrate that conditioning diffusion models on hierarchical cluster representations improves the generative performance on real-world datasets compared to other approaches. Moreover, a key strength of our method lies in its ability to generate images that are both representative and specific to each cluster, enabling more detailed visualization of the learned latent structure. Our approach addresses the generative limitations of VAE-based clustering approaches by leveraging their learned structure, thereby advancing the field of generative clustering.
This preprint has not undergone peer review or any post-submission improvements or corrections. The Version of Record of this contribution is published in Machine Learning and Knowledge Discovery in Databases. Research Track. ECML PKDD 2025, Lecture Notes in Computer Science (LNCS, vol. 16013), and is available online at https://doi.org/10.1007/978-3-032-05962-8_26.
@@ 1 Introduction
Generative modeling and clustering are two fundamental yet different tasks in machine learning. Generative modeling focuses on approximating the underlying data distribution, enabling the generation of new samples [14, 7]. Clustering, on the other hand, seeks to uncover meaningful and interpretable structures within data through the unsupervised detection of intrinsic relationships and dependencies [39, 6], facilitating better visualization and interpretation of the data. By integrating hierarchical dependencies into a deep latent variable model, TreeVAE [19] was recently proposed to bridge these two research directions. While TreeVAE is effective at hierarchical clustering, it falls short in generating high-quality images. Like other VAE-based models, it faces common issues such as producing blurry outputs [4]. In contrast, diffusion models [30, 11] have recently gained prominence for their superior image generation capabilities, progressively refining noisy inputs to produce sharp, realistic images.
Our work bridges this gap by introducing a second-stage diffusion model that is conditioned on the hierarchical cluster representations learned by TreeVAE. The proposed framework, TreeDiffusion, combines the strengths of both models to generate high-quality, cluster-specific images, achieving improved performance in image generation. The generative process begins by sampling the root embedding of a latent tree, which is learned during training. From there, the sample is propagated from the root to one leaf by (a) sampling a path through the tree and (b) applying a sequence of stochastic transformations to the root embedding along the chosen hierarchical path. Subsequently, the diffusion model harnesses the hierarchical information by conditioning its reverse diffusion process on the sampled path representations of the latent tree through a path encoder. A key strength of TreeDiffusion is its ability to generate images tailored to each cluster, providing enhanced visualization of the learned representations, as demonstrated by our qualitative results. For the same sample, our method can produce leaf-specific images that share common general properties but differ by features encoded in the latent hierarchy. Lastly, this approach overcomes the generative limitations of VAE-based hierarchical clustering models like TreeVAE while preserving their clustering performance.
1.1 Main Contributions
Our main contributions include (i) a unified framework that integrates hierarchical clustering into diffusion models, and (ii) a novel mechanism for controlling image synthesis based on learned clusters. We demonstrate that our approach (a) surpasses the generative limitations of VAE-based clustering models, and (b) produces samples that are both more representative of their respective clusters and closer to the true data distribution than models without hierarchical clustering integration.
@@ 2 Related Work
2.0.1 Variational Approaches for Hierarchical Clustering.
@@
Since their introduction, Variational Autoencoders (VAEs) [14] have been widely used for clustering tasks, due to their ability to learn structured latent representations [13]. For instance, one method integrates hierarchical Bayesian non-parametric priors to the latent space of VAEs by applying the nested Chinese Restaurant Processes to cluster the data based on infinitely deep and branching trees [8]. Other methods construct hierarchical representations during training, such as DeepECT [20] and TreeVAE [19]. While DeepECT aggregates data into a hierarchical tree in a single shared latent space, TreeVAE learns a tree-structured posterior distribution of latent stochastic variables. That is, TreeVAE models the data distribution by learning an optimal tree structure of latent variables. This results in latent embeddings that are automatically organized into a hierarchy, mimicking the hierarchical clustering process. Single-cell TreeVAE [35] further extends TreeVAE to single-cell RNA sequencing data by integrating batch correction, facilitating biologically plausible hierarchical structures. Despite their strong clustering performance, these models often fall short in generative quality, with few providing quantitative or qualitative evaluations of their sample generation capabilities.
2.0.2 Diffusion Models.
@@
Diffusion models have become state-of-the-art for image generation tasks over the past few years [30, 11, 31, 21, 5, 34, 27, 25]. One drawback of diffusion models is that their latent variables lack interpretability compared to the latent spaces of VAEs. To take advantage of the strengths of both approaches, researchers have explored architectures that combine the more interpretable latent spaces of VAEs with the advanced generative capabilities of diffusion models. Notable examples include DiffuseVAE [23], Diffusion Autoencoders [24], and InfoDiffusion [38]. Representation-conditioned image generation [17] illustrates how self-supervised learning can improve generative diffusion frameworks in unsupervised settings, reducing the gap between class-conditional and unconditional image generation.
2.0.3 Connecting Diffusion with Clustering.
@@
The research most closely related to our work focuses on using clustering as conditioning signals for diffusion models to improve generative quality. One approach [1] utilizes cluster assignments from k-means or TEMI clustering [2]. Similarly, another one [12] introduces a framework that employs the k-means clustering algorithm as an annotation function, generating self-annotated image-level, box-level, and pixel-level guidance signals. Both studies demonstrate the benefits of conditioning on clustering information to improve generative performance without going into the specifics of clustering performance itself. In contrast, our work further investigates which types of clustering information are most beneficial for the model, employing learned latent cluster representations alongside cluster assignments for conditioning. Related to conditioning on clusters, both kNN-Diffusion [28] and Retrieval-Augmented Diffusion Models [3] utilize nearest neighbor retrieval to condition generative models on similar embeddings, minimizing the need for large parametric models and paired datasets in tasks like text-to-image synthesis. Diffusion models have also been applied in incomplete multiview clustering to generate missing views to improve clustering performance [40, 41]. On a different note, recent research shows that training diffusion models is equivalent to solving a subspace clustering problem, explaining their ability to learn image distributions with few samples [37]. Finally, diffusion models have also been applied as a post-hoc method to enhance the generation quality of their multimodal clustering models [22]. However, to the best of our knowledge, no existing diffusion model explicitly uses hierarchical clustering to enhance the interpretability and generative performance of generative clustering models.
@@ 3 Method
We propose TreeDiffusion  [^1] The code is publicly available at https://github.com/JoGo175/TreeDiffusion., a two-stage framework consisting of a VAE-based generative hierarchical clustering model, followed by a hierarchy-conditional diffusion model. In the first stage, TreeVAE [19] serves as the clustering model, encoding hierarchical clusters within its latent tree structure, where the leaf nodes represent clusters. We select TreeVAE as it provides structured hierarchical latent representations from root to leaf, which are then processed by a path encoder to create the conditioning signal. In the second stage, a denoising diffusion implicit model (DDIM) [23], uses this conditioning signal to generate cluster-conditional samples. Hence, our model enables cluster-guided diffusion in unsupervised settings, analogously to classifier-guided diffusion [5] in supervised settings. Figure˜1 illustrates the workflow of TreeDiffusion. The following sections provide a detailed description of each stage of the model.
Figure 1: Schematic overview of the TreeDiffusion framework: TreeVAE encodes data into hierarchical latent variables, where a path is sampled from the root to a leaf node. An encoder network creates a conditioning signal using the sampled hierarchical path embeddings. The diffusion model leverages this information to condition its reverse process and generate a cluster-specific image.
@@ 3.1 Hierarchical Clustering with TreeVAE
The first part of TreeDiffusion involves a Tree Variational Autoencoder (TreeVAE) [19]. TreeVAE is a generative model that learns to hierarchically separate data into clusters through a latent tree structure. During training, the model dynamically grows a binary tree structure of stochastic variables. The process begins with a simple tree composed of a root and two child nodes, and it optimizes the corresponding ELBO over a fixed number of epochs. Afterward, the tree expands by adding two child nodes to an existing leaf node, prioritizing nodes with the highest assigned sample count to promote balanced leaves. This expansion continues iteratively, training only the subtree formed by the new leaves while freezing the rest of the model. This process repeats until the tree reaches a predefined depth or leaf count, alternating between optimizing model parameters and expanding the tree structure.
Let the set $\mathbb{V}$} represent the nodes of the tree. Each node corresponds to a stochastic latent variable, denoted as ${\mathbf{z}}_{0},\dots,{\mathbf{z}}_{V}$. Each latent variable follows a Gaussian distribution, whose parameters depend on their parent nodes through neural network layers. The set of leaves $\mathbb{L}$}, with ${\mathbb{L}}\subset{\mathbb{V}}$, represents the clusters present in the data. Starting from the root node, ${\mathbf{z}}_{0}$, a given sample traverses the tree to a leaf node, ${\mathbf{z}}_{l}$, in a probabilistic manner. The probabilities of moving to the left or right child at each internal node are determined by neural networks termed routers. These decisions, denoted by $\mathrm{c}_{i}$ for each non-leaf node $i$, follow a Bernoulli distribution, where $\mathrm{c}_{i}=0$ indicates the selection of the left child. The path $\mathcal{P}_{l}$ refers to the sequence of nodes from the root to one leaf $l$. Moreover, let $\bm{z}_{\mathcal{P}_{l}}=\left\{\bm{z}_{i}\mid i\in\mathcal{P}_{l}\right\}$ denote the set of latent embeddings for each node in the path $\mathcal{P}_{l}$. The latent tree encodes a sample-specific probability distribution of paths. Each leaf embedding, ${\mathbf{z}}_{l}$ for $l\in{\mathbb{L}}$, represents the learned data representations. In TreeVAE, leaf-specific decoders use these embeddings to reconstruct or generate new cluster-specific images, i.e. given a dataset $\bm{X}$}, TreeVAE reconstructs $\hat{{\bm{X}}}=\{\hat{{\bm{X}}}^{(l)}\mid l\in{\mathbb{L}}\}$. The generative model (1) and inference model (2) of TreeVAE are defined as follows, according to [19]:
@@
3.2 Diffusion conditioned on Hierarchical Clusters
The second part of TreeDiffusion incorporates a conditional diffusion model. We assume the same forward process as in standard Denoising Diffusion Probabilistic Models (DDPM) [11], which gradually introduces noise to the data $\bm{x}_{0}$ over $T$ steps. The intermediate states, $\bm{x}_{t}$ for $t=1,\dots,T$, follow a trajectory determined by a noise schedule $\beta_{1},\dots,\beta_{T}$} that controls the rate of data degradation, whereby $\alpha_{t}=\left(1-\beta_{t}\right)$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. Hence, the forward process can be summarized as follows:
@@ 3.2 Diffusion conditioned on Hierarchical Clusters
$q\left(\bm{x}_{t}\mid\bm{x}_{0}\right)$ $=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},\left(1-\overline{\alpha}_{t}\right)\bm{I}\right)$ (5)
For the reverse process, TreeDiffusion starts with random noise, similar to standard diffusion models. TreeDiffusion relies exclusively on the latent hierarchical information provided by TreeVAE, which is based on the tree structure learned during the first stage. Specifically, a tree leaf $l$ corresponds to a selected cluster, determined by sampling from the TreeVAE path probabilities. Thus, this leaf represents a unique path through the hierarchical structure. The hierarchical conditioning information is derived from $\bm{z}_{\mathcal{P}_{l}}$, the set of latent embeddings along the path from the root node to the chosen leaf. These embeddings are further processed by a dedicated path encoder, which aggregates the information to produce the conditioning signal ${\bm{y}}_{l}$:
${\bm{y}}_{l}=\sum_{i\in\mathcal{P}_{l}}\left(f_{\text{embed}}(\bm{z}_{i})+f_{\text{node}}(i)\right).$
Here, $f_{\text{embed}}$ and $f_{\text{node}}$ are each implemented as projection blocks consisting of two MLP layers with an activation in-between, and jointly trained with the diffusion model. For each node in the path, its embedding and corresponding node index are projected independently into the time embedding dimension of the U-Net decoder [26, 21]. The resulting projections are aggregated into the unified conditioning signal ${\bm{y}}_{l}$, which is then combined with the time-step embeddings to guide the U-Net during the denoising process. Consequently, this conditioning mechanism directly influences the reverse process. Let $\psi$ denote the parameters of the denoising model, and let $p(l|{\bm{x}}_{0})$ be the probability that the sample ${\bm{x}}_{0}$ is assigned to leaf $l$ in the latent tree. The reverse process can then be summarized as follows:
[…]
@@
The path sampling ensures that different leaves are considered, encouraging the diffusion model to perform effectively across all leaves. Consequently, our approach addresses the distinct clusters inherent to TreeVAE, allowing the model to adapt and encouraging cluster-aware refinements in the images. This guidance in the image generation process assists the denoising model in learning cluster-specific image reconstructions.
The following design considerations are implemented in TreeDiffusion to achieve computational efficiency without compromising effectiveness. Due to the large number of denoising steps required, DDPM sampling can be computationally expensive. To address this issue, we opt for the DDIM sampling procedure [31] instead of the standard DDPM [11]. DDIMs significantly accelerate inference by using only a subset of denoising steps, making the process more efficient while maintaining high-quality results. Finally, by employing a two-stage training strategy, where the conditional diffusion model is trained using a pre-trained TreeVAE model, TreeDiffusion preserves the hierarchical clustering performance of TreeVAE. Hence, we can combine the effective clustering of TreeVAE with the superior image generation capabilities of diffusion models.